OpenStudy (anonymous):

What would the equation of the line normal to this surface be? x^2-y^2+z^2=4 at the point (2,-3,3)

OpenStudy (anonymous):

I know the vector is <2x,-2y,2z> but how do I do the equation? @amistre64

OpenStudy (amistre64):

how do we define a line?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (amistre64):

no ... you are working with vectors here. think parametrically

OpenStudy (anonymous):

2xi-2yj+2zk?

OpenStudy (amistre64):

given a point (xo,yo,zo) and a direction vector (a,b,c) the parametric form of a line is defined as: x = xo + at y = yo + bt z = zo + ct

OpenStudy (amistre64):

your gradient can be 'simplified' since the length of it is immaterial: factor our the 2 :)

OpenStudy (anonymous):

oh I didn't know I could do that thank you

OpenStudy (amistre64):

a direction is whats important to: <2x,-2y, 2z> is in the same direction as: <x,-y,z> right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so then x=2+xt, y=-3-yt, z=3+zt ?

OpenStudy (amistre64):

well, almost. you need to define the gradient by the point as well

OpenStudy (amistre64):

<2x,-2y,2z> @ (2,-3,3) = <2(2),-2(-3),2(3)> = 2<2,3,3> x = 2+2t y = -3+3t z = 3+3t

OpenStudy (anonymous):

ohh i get it and so those are the equations of the normal line?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ah got it and what if i wanted to do the equation for the tangent plane?

OpenStudy (amistre64):

the gradient tells us the normal at a given point the line therefore is just the parametricized setup

OpenStudy (amistre64):

we have a normal, dot it to all the vectors (x,y,z) to (2,-3,3)

OpenStudy (amistre64):

when a dot product is equal to 0, then the vectors are perpendicular

OpenStudy (anonymous):

gotcha thank youuu!! :)

OpenStudy (amistre64):

youre welcome all vectors from (2,-3,3) to (x,y,z) = (x-2, y+3, z-3) dot this with our gradient to get (2,3,3).(x-2, y+3, z-3)=0 2(x-2)+3(y+3)+3(z-3)=0 simplify as wanted